Q.let s be a square of unit area.consider a quadilateral which has 1 vertex on each side of the square ,if a,b,c,d denote the length of the side of the quadilateral ...P.T. 2<a^1+b^2+c^2+d^2<4

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Aman,

Ans:- Let the square be ABCD and the corresponding points be PQRS( P on AB Q on BC etc)

then we get the following eq

BP^2+BQ^2=a^2

AP^2+AS^2=b^2

SD^2+DR^2=c^2

QC^2+CR^2=d^2

Hence adding we get,

AP^2+BP^2+BQ^2+QC^2+CR^2+RD^2+SD^2+AS^2=a^2+b^2+c^2+d^2 =X(let).................(1)

breaking eq (1) once in a^2+b^2=(a+b)^2 - 2ab form and again (a-b)^2+2ab form we get

X=(AP+BP)^2+(BQ+QC)^2+(CR+RD)^2+(SD+AS)^2 -2∑(AP.PB)

=4 - 2∑(AP PB)

X<4

AGAIN

X=∑(AP - PB)^2 +2∑(AP PB)

Hence X>2

So WE GET

2<X<4(Proved)

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.
win exciting gifts by answering the questions on Discussion Forum.

All the best Aman !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR